OpenStudy (anonymous):

Determine whether the following lines are parallel, perpendicular, or the same line. y = 3x – 4 and y – 3 = 3( x + 1) A. Parallel B. Perpendicular C. Same Line D. None of these

OpenStudy (jdoe0001):

solve the 2nd for "y", what do you get?

OpenStudy (anonymous):

y = 3x – 4 --------- I y – 3 = 3( x + 1) y - 3 = 3x + 3 y = 3x + 3 + 3 y = 3x + 6 --------- II Since both I and II have same slope but different y-intercepts, so they will be parallel lines